When 3x2 - 8x + 5 is factored, one of the factors is:
x - 1
3x + 5
x - 5
3x - 1
x + 5

Answer by checkley71(8403)

3x^2-8x+5=0
(3x-5)(x-1)=0
(x-1) is a factor